OpenStudy (anonymous):

In a certain year the amount A of garbage in pounds produced after t days by an average household is given by A=5.5t Graph the equation t>0 How many days it take for the average household 100 pounds of garbage?

OpenStudy (anonymous):

amount A of garbage in pounds produced after t days A = 5.5 t So, it takes t days to produce amount A of garbage, => t = A / 5.5 Substitute A = 100 in above equation to find the number of days takes to produces 100 pounds of garbage You get t = 100/5.5 

OpenStudy (anonymous):

I am veiwning it

OpenStudy (anonymous):

veiwing

OpenStudy (anonymous):

do i put 100 divide into 5.5 to figure out the approximate amount of use in 100days

OpenStudy (anonymous):

yes, you should get t as: 100/5.5 = 18.18 days

OpenStudy (anonymous):

yeah i have that

OpenStudy (anonymous):

i gotta round to the nearest tenth

OpenStudy (anonymous):

18.2

OpenStudy (anonymous):

@sistergrl4 Great! That's the correct answer:)

OpenStudy (anonymous):

im sorry it says to nearest whole number

OpenStudy (anonymous):

Then the answer is 18!

OpenStudy (anonymous):

but we have 18.18 round to 18.2

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

do you have time for another question?

OpenStudy (anonymous):

Yes of course:)

OpenStudy (anonymous):

ok here it goes

OpenStudy (anonymous):

A certain mountain has an elevation of 19,809 feet. In 1917, the glacier covered 4 acres. By 2005 this glacier has melted 1 acre. The year change is blank acres/year

OpenStudy (anonymous):

What is the equation that gives the acreage of the glacier t years past 1917. A= blank

OpenStudy (anonymous):

I have A=-3/88

OpenStudy (anonymous):

this problem is difficult

OpenStudy (anonymous):

Let's assume this glacier melts at a constant rate each year. We have to find the equation of the acreage. Here, we have two point for (1917,4), (2005,1) Let slope of the equation be A A = (y1 - y2) / (x1 - x2) = (1 - 4) / (2005- 1917) = -3/88 the slope states that 3 acres are lost every 88 years which works out to -1/29.333 which is one acre every 29.333 years or 0.0340909...acres per year Now you have found A, Ok the equation that gives the acreage of the glacier t years past 1917: x = At + b Then in 1917, t =0 4 = -3/88 * 0 + b => b =4 x = -3/88t -4 is the equation where a= acres of glacier remaining after t years.

OpenStudy (anonymous):

ok so how would i wright acres of change over years is it -3/88

OpenStudy (anonymous):

Sorry this shoud be x = -3/88t +4

OpenStudy (anonymous):

so this should be A= -3/88+4

OpenStudy (anonymous):

it also ask me to fill in the blank for acres/years?

OpenStudy (anonymous):

yes, this is the answer to your problem, A= -3/88 t +4

OpenStudy (anonymous):

The yearly rate change blank is acres/years?

OpenStudy (anonymous):

ok I have the A= answer right for the -3/88...

OpenStudy (anonymous):

The yearly rate change is 3/88

OpenStudy (anonymous):

ok that is what i am assuming because i was confused and also do i still have to place the negative sign on the fraction or no?

OpenStudy (anonymous):

I really do appreciate you for taking the time out to help me

OpenStudy (anonymous):

yes, it's a negative rate, so you have to put the - sign: -3/88

OpenStudy (anonymous):

ok do you have time for one last question this is my last problem i worked it out but idw if i did it right

OpenStudy (anonymous):

@sistergrl4 It's a pleasure:) yes, just ask me:)

OpenStudy (anonymous):

f(a)=t,f(b)=s,f(c)=q,f(d)=t Write the function set pair for domain and range

OpenStudy (anonymous):

I have f= (a,t), (b,s),(c,q),(d,t)

OpenStudy (anonymous):

domain = a,b,c,d range= t,s,q,t

OpenStudy (anonymous):

Good! the domain is {a,b,c,d} and the Range = {t,s,q}

OpenStudy (anonymous):

ok i was asking because usually this type of prob has numbers not alphabets

OpenStudy (anonymous):

do you like coming on here a lot to help people i am not a superb math student but i try at my best

OpenStudy (anonymous):

Yes, functions can also be extended to sets of ordered pairs [see the graph]  Yes, I'll do my best to help u whenever I'm free! I know you'll soon become a superb mathematician! All the best!!:)

OpenStudy (anonymous):

yeah thanks very much